### v1

# Uncertainty relations for overcomplete measurements from generalized equiangular tight frames

Abstract
The current study aims to examine uncertainty relations for overcomplete measurements from generalized equiangular tight frames. Informationally overcomplete measurements are a valuable tool in quantum information processing, including tomography and state estimation. The maximal sets of mutually unbiased bases are the most common case of informationally overcomplete measurements. The existence of $d+1$ of mutually unbiased bases is proved for $d$ being a prime power. More general classes of informationally overcomplete measurements have been proposed for various purposes. Such measurements are typically characterized by some inner structure maintaining the required properties. It leads to restrictions imposed on generated probabilities. To apply the measurements of interest, these restrictions should be converted into information-theoretic terms. To describe the amount of uncertainty quantitatively, we use the Tsallis and Rényi entropies as well as probabilities of separate outcomes. The obtained results are based on estimation of the index of coincidence. The derived relations are briefly exemplified.
overcomplete measurements, index of coincidence, Tsallis entropy, Rényi entropy, conical designs
I Introduction
Quantum technologies of information processing are currently the subject of intensive studies [1], though complete awareness of their role belongs to the future. The final stage of various protocols for quantum information processing involves measurements required for obtaining information about resulting quantum states. Quantum measurements are represented by positive operator-valued measures (POVMs) [2]. Each of them consists of positive semidefinite operators that sum up to the identity operator. To ensure convenient dealing with, the utilized measurements are built to have good properties. A measurement is informationally complete, when every possible state is uniquely determined by the measurement statistics [3]. Complete sets of mutually unbiased bases are very important example [4, 5, 6]. Symmetric informationally complete measurements (SIC-POVMs) have many nice properties to use [7, 8]. Fisher-symmetric informationally complete measurements and their applications were discussed in [9, 10]. Experimental implementation of symmetric informationally complete measurements were described in [11, 12, 13, 14, 15].
In fact, the problem of building desired structures in Hilbert spaces is often hard to resolve [8, 16]. It is known that $d+1$ of mutually unbiased bases in dimension $d$ exist for $d$ being a prime power. But the maximal number of mutually unbiased bases remains unknown even for $d=6$ [17]. There are reasons to believe that SIC-POVMs exist for all $d$, and this conjecture is due to Zauner [18]. Here, there are ways to weak some of the imposed restrictions. Dealing with rank-one measurement operators, we may reduce the number of outcomes. In this regard, equiangular tight frames are shown to deserve more attention than they have obtained [19]. Equiangular measurements are a special case of more general concepts such as equioverlapping measurements [20, 21] and the so-called $(1,M)$-POVMs [22]. Another way is to allow measurement operators of arbitrary rank. General symmetric informationally complete measurements were built [23, 24] and applied [25]. Mutually unbiased measurements instead of bases were proposed in [26]. Recently, generalized symmetric measurements [27] and generalized equiangular ones [28] have been studied.
The Heisenberg uncertainty principle [29] is widely recognized as a fundamental concept that the observer perturbs what he observes. Since the first formal derivations of Kennard [30] and Robertson [31] appeared, many approaches and scenarios were addressed [32]. For finite-dimensional systems, entropic uncertainty relations [33, 34] give a useful alternative to Robertson’s formulation. The well-known Maassen–Uffink uncertainty relation [34] was originally conjectured by Kraus [35]. This relation becomes a tool in building feasible schemes to detect of nonclassical correlations such as entanglement [36, 37] and steerability [38, 39, 40]. Entropic uncertainty relations have also emerged in the security analysis of quantum cryptographic protocols [41, 42, 43]. Entropic formulation gives a natural way to address uncertainties in the presence of quantum memory [44]. For more results on entropic uncertainty relations, see the reviews [45, 46, 47] and references therein. For measurements with a special inner structure, the Maassen–Uffink uncertainty relation can be improved [48, 49, 50, 51].
It is important to examine properties of specially built quantum measurements as completely as possible, even if the final opinion on their use is now unformed. The aim of this study is to derive uncertainty relations for overcomplete measurements from generalized equiangular tight frames. It immediately concerns uncertainty relations for generalized symmetric measurements due to the correspondence mentioned in [28]. The new relations are expressed in terms of the Renyi [52] and Tsallis entropies [53] as well as sums of probabilities. Some of the used inequalities are based on considering information diagrams as described in [54, 55]. The paper is organized as follows. Section II introduces the preliminary material concerning generalized equiangular measurements and generalized symmetric ones. Some properties of the Tsallis and Rényi entropies are recalled as well. New uncertainty relations for the considered measurements are presented in Sec. III. A utility of the obtained inequalities is briefly exemplified. Section IV concludes the paper with a summary of the results.
II Preliminaries
This section aims to recall the required material concerning informationally overcomplete measurements derived from generalized equiangular tight frames. It also considers entropies that will be used to characterize uncertainties in quantum measurements of interest.
II.1 Generalized equiangular measurements and conical $2$-designs
Basic reasons to put informationally overcomplete measurements are discussed in [28]. We recall only very necessary facts about them. Let ${\mathcal{H}}_{d}$ be $d$-dimensional Hilbert space. By ${\mathcal{L}}({\mathcal{H}}_{d})$ and ${\mathcal{L}}_{+}({\mathcal{H}}_{d})$, we respectively mean the space of linear operators on ${\mathcal{H}}_{d}$ and the set of positive semidefinite ones. The state of a quantum system is described by density matrix $\boldsymbol{\rho}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ normalized as $\mathrm{tr}(\boldsymbol{\rho})=1$. In general, any quantum measurement is represented as a set =${\mathcal{E}}=\bigl{\{}{\mathsf{E}}_{j}\bigr{\}}_{j=1}^{N}$ with ${\mathsf{E}}_{j}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ such that the completeness relation holds,
[…]
For the pre-measurement state $\boldsymbol{\rho}$, the probability of $j$-th outcome is equal to $\mathrm{tr}({\mathsf{E}}_{j}\boldsymbol{\rho})$. In principle, the number of different outcomes can exceed the dimensionality. This property is indispensable in many questions of quantum information science.
A collection of $M$ generalized equiangular lines ${\mathcal{P}}_{\mu}=\bigl{\{}{\mathsf{P}}_{\mu,j}\bigr{\}}_{j=1}^{N_{\mu}}$ with ${\mathsf{P}}_{\mu,j}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ is a generalized equiangular measurement if the following items hold [28]:
[…]
$\sum_{\mu=1}^{M}\gamma_{\mu}$ $=1\,,$ (3)
$\mathrm{tr}({\mathsf{P}}_{\mu,j})$ $=a_{\mu}\,,$ (4)
$\mathrm{tr}({\mathsf{P}}_{\mu,j}^{2})$ $=b_{\mu}\mathrm{tr}({\mathsf{P}}_{\mu,j})^{2}\,,$ (5)
$\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{P}}_{\mu,j})$ $=c_{\mu}\mathrm{tr}({\mathsf{P}}_{\mu,i})\,\mathrm{tr}({\mathsf{P}}_{\mu,j})\qquad(i\neq{j})\,,$ (6)
$\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{P}}_{\nu,j})$ $=f\,\mathrm{tr}({\mathsf{P}}_{\mu,j})\,\mathrm{tr}({\mathsf{P}}_{\nu,j})\qquad(\mu\neq\nu)\,.$ (7)
From the above definition, it follows that $f=d^{-1}$,
$a_{\mu}=\frac{\gamma_{\mu}d}{N_{\mu}}\ ,\qquad c_{\mu}=\frac{N_{\mu}f-b_{\mu}}{N_{\mu}-1}\ .$ (8)
The admissible range for $b_{\mu}$ is obtained as [28]
$\frac{1}{d}<b_{\mu}\leq\frac{\min\{d,N_{\mu}\}}{d}\ .$ (9)
A single equiangular measurement is a particular case for $M=\gamma=b=1$. It should be noted here that our notation slightly differs from the notation of [28]. The generalized equiangular measurement can also be interpreted as a set of $M$ POVMs so that $\mu$-th POVM has elements $\gamma_{\mu}^{-1}\,{\mathsf{P}}_{\mu,j}$. Due to this correspondence, a generalized equiangular measurement relates to some generalized symmetric measurement [28].
The generalized symmetric measurements were introduced as collections of POVMs that are not equinumerous [27]. It is important here that elements of generalized equiangular measurements can be constructed from elements of generalized symmetric measurements. One uses [27] a set of $M$ POVMs ${\mathcal{E}}_{\mu}=\bigl{\{}{\mathsf{E}}_{\mu,j}\bigr{\}}_{j=1}^{N_{\mu}}$ satisfying additional symmetry conditions:
$\mathrm{tr}({\mathsf{E}}_{\mu,j})$ $=w_{\mu}\,,$ (10)
$\mathrm{tr}({\mathsf{E}}_{\mu,j}^{2})$ $=x_{\mu}\,,$ (11)
$\mathrm{tr}({\mathsf{E}}_{\mu,i}{\mathsf{E}}_{\mu,j})$ $=y_{\mu}\qquad(i\neq{j})\,,$ (12)
$\mathrm{tr}({\mathsf{E}}_{\mu,i}{\mathsf{E}}_{\nu,j})$ $=z_{\mu\nu}\qquad(\mu\neq\nu)\,.$ (13)
For informationally complete measurements, one has
$\sum_{\mu=1}^{M}N_{\mu}=d^{2}+M-1\,.$ (14)
There are four construction methods with the use of linear combinations of Hermitian orthonormal operator bases [27]. We refrain from presenting the details here.
The following observation has been made in [28]. Suppose that positive numbers $\gamma_{\mu}$ obey (3) and $M$ POVMs ${\mathcal{E}}_{\mu}=\bigl{\{}{\mathsf{E}}_{\mu,j}\bigr{\}}_{j=1}^{N_{\mu}}$ satisfy (10)–(13). Then operators of the form
${\mathsf{P}}_{\mu,j}=\gamma_{\mu}\,{\mathsf{E}}_{\mu,j}$ (15)
compose a generalized equiangular measurement. The correspondence between the coefficients reads as
$a_{\mu}=\gamma_{\mu}w_{\mu}\,,\qquad b_{\mu}=\frac{x_{\mu}}{w_{\mu}^{2}}\ ,\qquad c_{\mu}=\frac{y_{\mu}}{w_{\mu}^{2}}\ ,\qquad f=\frac{z_{\mu\nu}}{w_{\mu}w_{\nu}}\ .$ (16)
The obtained POVM is informationally complete, if it is constructed from informationally complete set of elements ${\mathsf{E}}_{\mu,j}$ and $\gamma_{\mu}>0$ for all $\mu=1,\ldots,M$ [28]. Thus, the generalized symmetric measurements are sufficient to build a generalized equiangular measurement. Despite of simplicity of the relation (15), this construction has some nice features. For example, it is possible to manipulate the coefficients in such a way that both the traces $\mathrm{tr}({\mathsf{P}}_{\mu,j}^{2})$ and $\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{P}}_{\mu,j})$ become $\mu$-independent [28].
Projective designs have found use in various questions of quantum information theory. It seems that the special case of $2$-designs is provided with a large number of examples. Let $\bigl{\{}|\phi_{j}\rangle\bigr{\}}_{j=1}^{N}$ be a set of $N$ unit vectors of ${\mathcal{H}}_{d}$. There are several equivalent definitions, when these vectors form a projective $2$-design [18, 56]. One of them reads as
$\frac{1}{N}\,\sum_{j=1}^{N}|\phi_{j}\rangle\langle\phi_{j}|\otimes|\phi_{j}\rangle\langle\phi_{j}|=\frac{2}{d^{2}+d}\>{\mathsf{\Pi}}_{\mathrm{sym}}\,,$ (17)
where ${\mathsf{\Pi}}_{\mathrm{sym}}$ is the orthogonal projection on the symmetric subspace of ${\mathcal{H}}_{d}\otimes{\mathcal{H}}_{d}$. Conical designs are generalizations of the above concept to positive operators [57, 58]. We introduce the unitary swap operator $\tilde{\mathsf{W}}$ on ${\mathcal{H}}_{d}\otimes{\mathcal{H}}_{d}$ which takes $|\psi\rangle\otimes|\varphi\rangle$ to $|\varphi\rangle\otimes|\psi\rangle$. The set $\bigl{\{}{\mathsf{A}}_{j}\bigr{\}}_{j=1}^{N}$ with ${\mathsf{A}}_{j}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ is a conical $2$-design, if and only if
[…]
where $\kappa_{+}\geq\kappa_{-}>0$. It can be noticed that $2\kappa_{\pm}=\kappa_{\mathrm{s}}\pm\kappa_{\mathrm{a}}$. Taking the partial trace in (18) yields
[…]
To each conical $2$-design, we can assign POVM =${\mathcal{E}}=\bigl{\{}{\mathsf{E}}_{j}\bigr{\}}_{j=1}^{N}$ with elements
${\mathsf{E}}_{j}=\frac{\mathrm{tr}({\mathsf{A}}_{j})}{\kappa_{+}d+\kappa_{-}}\,{\mathsf{A}}_{j}\,.$ (20)
The case with constant $\mathrm{tr}({\mathsf{A}}_{j})$ is very close to a projective $2$-design. On the other hand, a possibility to manipulate the coefficients generally excludes this restriction.
The author of [28] showed that some of the generalized equiangular measurements are conical $2$-designs. Let us take the coefficients such that
$a_{\mu}^{2}(b_{\mu}-c_{\mu})=S$ (21)
for all $\mu$. Then the generalized equiangular measurement is a conical 2-design with
$\kappa_{+}=\frac{\sigma-S}{d}\,,\qquad\kappa_{-}=S\,,\qquad\sigma=\sum_{\mu=1}^{M}a_{\mu}\gamma_{\mu}\,.$ (22)
There exist conical $2$-designs even though they are constructed from the generalized symmetric measurements that are not conical $2$-designs [28]. This property could be useful in producing new conical $2$-designs. Unitary designs are now considered as a powerful tool in quantum information science [59, 60]. Thus, generalized equiangular measurements deserve to be studied in application to questions of quantum information science. It turned out that generated probabilities satisfy certain conditions that lead to various restrictions including uncertainty relations. To express the amount of uncertainties, the Tsallis and Rényi entropies will be used in this paper.
II.2 Generalized entropies and the index of coincidence
The index of coincidence is used in various questions of information theory [61]. Let $P=(p_{1},\ldots,p_{N})$ be a probability distribution with the index running over the set $\{1,\ldots,N\}$. The corresponding index of coincidence reads as
$I(P)=\sum_{j=1}^{N}p_{j}^{2}\,.$ (23)
Indicies with larger degrees of probabilities have also found use [61]. Applications of such indicies to derive uncertainty relations were given in [62, 63, 64]. For $\alpha>0$, the Tsallis $\alpha$-entropy is expressed as [53]
$H_{\alpha}(P)=\frac{1}{1-\alpha}\,\biggl{(}\,\sum_{j=1}^{N}p_{j}^{\alpha}-1\biggr{)}=\sum_{j=1}^{N}p_{j}\ln_{\alpha}\biggl{(}\frac{1}{p_{j}}\biggr{)}.$ (24)
The right-hand side of (24) uses the $\alpha$-logarithm of positive variable,
$\ln_{\alpha}(X)=\begin{cases}\frac{X^{1-\alpha}-1}{1-\alpha}\>,&\text{if}\ 0<\alpha\neq 1\,,\\ \ln{X}\,,&\text{if}\ \alpha=1\,.\end{cases}$ (25)
Substituting the probabilities $\mathrm{tr}({\mathsf{E}}_{j}\boldsymbol{\rho})$ into (24) gives the entropy $H_{\alpha}({\mathcal{E}};\boldsymbol{\rho})$. Some results of the current study deal with the Rényi $\alpha$-entropy defined as [52]
$R_{\alpha}(P)=\frac{1}{1-\alpha}\,\ln\biggl{(}\,\sum_{j=1}^{N}p_{j}^{\alpha}\biggr{)}\,.$ (26)
In the limit $\alpha\to\infty$, the right-hand side of (26) leads to the so-called min-entropy equal to the minus logarithm of the maximal probability. For $\alpha=1$, both the entropies (24) and (26) reduce to the Shannon entropy,
$H_{1}(P)=-\sum_{j=1}^{N}p_{j}\ln{p}_{j}\,.$ (27)
Due to the obvious connection
$R_{\alpha}(P)=\frac{1}{1-\alpha}\,\ln\bigl{\{}1+(1-\alpha)H_{\alpha}(P)\bigr{\}}\,,$ (28)
each of Tsallis-entropy inequalities may lead to a Rényi-entropy counterpart. In more detail, applications of the Tsallis and Rényi entropies in quantum physics are described in [65].
For the given index of coincidence and $\alpha\in(0,2]$, we have
$H_{\alpha}(P)\geq\ln_{\alpha}\biggl{(}\frac{1}{I(P)}\biggr{)}.$ (29)
This fact follows from Jensen’s inequality. The paper [54] showed the way to improve (29) due to information diagrams. To each probability distribution, we assign the index of coincidence and the entropy of interest. Using the former as the abscissa and the latter as the ordinate, each probability distribution is shown by a point in the plane [61]. The problem consists in describing the boundaries of the resulting set of points. In effect, the answer depends on the actual number of nonzero probabilities. It is known that the lower-bounding curve passes through points of the form $\bigl{(}k^{-1},\ln_{\alpha}(k)\bigr{)}$ with integer $k\geq 1$. It corresponds to the uniform probability distribution with $k$ elements.
It was demonstrated for information diagrams with the Shannon entropy that the theoretically best lower-bounding curve is not smooth [61]. Expectedly, the same feature takes place for the Tsallis $\alpha$-entropy. It can be shown that a smoothness is violated in the points with abscissas $X=k^{-1}$. In more detail, this question is explained in [54, 55]. Actually, the inequality (29) is improved by replacing the graph of smooth function $X\mapsto\ln_{\alpha}\bigl{(}X^{-1}\bigr{)}$ with the polygonal line connecting the points with abscissas $X=k^{-1}$ and ordinates $\ln_{\alpha}(k)$ for integer $k\geq 1$. New graph shows the piecewise linear function $X\mapsto{L}_{\alpha}(X)$ expressed as [54]
$L_{\alpha}(X)=u_{\alpha{k}}-v_{\alpha{k}}\,X\,,\qquad X\in\biggl{[}\frac{1}{k+1}\,,\frac{1}{k}\biggr{]}\,.$ (30)
Here, the coefficients $u_{\alpha{k}}$ and $v_{\alpha{k}}$ read as
$u_{\alpha{k}}$ $=(k+1)\ln_{\alpha}(k+1)-k\ln_{\alpha}(k)\,,$ (31)
$v_{\alpha{k}}$ $=k(k+1)\bigl{\{}\ln_{\alpha}(k+1)-\ln_{\alpha}(k)\bigr{\}}\,.$ (32)
Due to the form of the polygonal line, the function $X\mapsto{L}_{\alpha}(X)$ is decreasing and convex. It then holds for $\alpha\in(0,2]$ that [54]
$H_{\alpha}(P)\geq{L}_{\alpha}\bigl{(}I(P)\bigr{)}\,.$ (33)
This inequality is not tight, except for points of the form $\bigl{(}k^{-1},\ln_{\alpha}(k)\bigr{)}$. For $\alpha=1$, the inequality (33) reduces to one of the main results of [61]. We will apply (33) to derive uncertainty relations for informationally overcomplete measurements and conical $2$-designs.
Information diagrams have also been applied to the maximal probability. Each diagram now fills so that the ordinate shows values of the maximal probability [54]. The upper bound on the maximal probability at the given index of coincidence was presented in [48]. Let $X\mapsto\varLambda_{p}(X)$ be a piecewise smooth function defined as
$\varLambda_{p}(X)=\frac{1}{k}\,\biggl{(}1+\sqrt{\frac{kX-1}{k-1}}\>\biggr{)}\,,\qquad X\in\biggl{[}\frac{1}{k}\,,\frac{1}{k-1}\biggr{]}\,.$ (34)
This function describes the boundary of the corresponding information diagram from below [54]. Hence, we obtain a two-sided estimate of the maximal probability, viz.
$\varLambda_{p}\bigl{(}I(P)\bigr{)}\leq\underset{1\leq{j}\leq{N}}{\max}\,p_{j}\leq\frac{1}{N}\,\Bigl{(}1+\sqrt{N-1}\,\sqrt{NI(P)-1}\,\Bigr{)}\,,$ (35)
where $N$ is the number of possible outcomes. The left-hand side of (35) is stronger than an obvious estimate $I(P)\leq\max\,p_{j}$. The latter is tight only in points of the form $\bigl{(}k^{-1},k^{-1}\bigr{)}$ with integer $k\geq 1$. Each of these points of the diagram corresponds to a uniform probability distribution [54, 55].
One of the results of [55] allows one to estimate sums of two probabilities from above at the given index of coincidence. Let the index of coincidence of the distribution $P=(p_{1},\ldots,p_{N})$ obey $N^{-1}\leq{I}(P)\leq 2^{-1}$. It then holds that
$\underset{i\neq{j}}{\max}\,\bigl{\{}p_{i}+p_{j}\bigr{\}}\leq\frac{1}{N}\,\Bigl{(}2+\sqrt{2N-4}\,\sqrt{NI(P)-1}\,\Bigr{)}\,.$ (36)
The inequality (36) is formally correct for all acceptable values of $I(P)$, but a nontrivial bound is obtained only for $I(P)<2^{-1}$. As a rule, the number of outcomes in overcomplete measurements is sufficiently large. Hence, the last inequality is typically satisfied. For example, this holds for a complete set of $d+1$ MUBs in dimension $d\geq 3$ [55].
III Main results
This section presents uncertainty relations for informationally overcomplete measurements generalized equiangular tight frames including conical $2$-designs. One shall address the entropic formulation as well as the case of probabilities for separate outcomes. All the derived relations will be exemplified.
III.1 Entropic uncertainty relations
Informational entropies [33, 34] provide a flexible way to pose the uncertainty principle for finite-dimensional systems. In some respects, the entropic formulation have advantages in comparison with the traditional one [31]. Entropic uncertainty relations are important not only from the conceptual viewpoint, but also for a lot of applications [45, 46]. In particular, such relations are used in derivation of criteria to detect entanglement [36, 37] and steerability [38, 39, 40]. Statistics of the measurements of interest can be interpreted in two ways. First, it relates to a single generalized equiangular measurement. Second, one links to a set of generalized symmetric measurements. For these measurements [28], the involved sums of squared probabilities satisfy a certain relation. This fact leads to a lot of consequences. To avoid bulky expressions, we will use positive numbers
$\omega_{\mu}$ $=\frac{1}{\mathcal{C}}\>\frac{\gamma_{\mu}^{2}}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}\ ,$ (37)
$\mathcal{C}$ $=\sum_{\mu=1}^{M}\frac{\gamma_{\mu}^{2}}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}\ .$ (38)
By construction, one has
$\sum_{\mu=1}^{M}\omega_{\mu}=1\,.$ (39)
Thus, the number (37) can be treated as a weight related to the $\mu$-th equiangular line. The following statement takes place.
Proposition 1
Let ${\mathbb{E}}=\bigl{\{}{\mathcal{E}}_{\mu}\bigr{\}}$ be a set of $M$ generalized symmetric POVMs assigned to a generalized equiangular measurement due to (15). For $\alpha\in(0,2]$, we have
$\sum_{\mu=1}^{M}\omega_{\mu}\,H_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})\geq{L}_{\alpha}\biggl{(}\frac{d\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}\gamma_{\mu}^{-1}a_{\mu}\omega_{\mu}\biggr{)}\,.$ (40)
Let ${\mathcal{P}}=\bigcup\,{\mathcal{P}}_{\mu}$ be a generalized equiangular measurement whose elements form a conical $2$-design with the parameters (22). For $\alpha\in(0,2]$, it holds that
$H_{\alpha}({\mathcal{P}};\boldsymbol{\rho})\geq{L}_{\alpha}\bigl{(}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+(\sigma-S)d^{-1}\bigr{)}\,.$ (41)
Proof. It was shown in appendix C of [28] that
$\mathrm{tr}(\boldsymbol{\rho}^{2})-\frac{1}{d}=\sum_{\mu=1}^{M}\frac{1}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}\biggl{(}\sum_{j=1}^{N_{\mu}}\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})^{2}-\frac{a_{\mu}\gamma_{\mu}}{d}\biggr{)}.$ (42)
Combining this with (15) and (37) immediately gives
$\mathcal{C}\,\sum_{\mu=1}^{M}\omega_{\mu}\,I_{\mu}=\mathrm{tr}(\boldsymbol{\rho}^{2})-\frac{1}{d}+\frac{1}{d}\,\sum_{\mu=1}^{M}\frac{\gamma_{\mu}}{a_{\mu}(b_{\mu}-c_{\mu})}\ ,$ (43)
where $I_{\mu}\equiv{I}({\mathcal{E}}_{\mu};\boldsymbol{\rho})$. Using (39), we further write
$\sum_{\mu=1}^{M}\omega_{\mu}\,H_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})$ $\geq\sum_{\mu=1}^{M}\omega_{\mu}\,L_{\alpha}(I_{\mu})$ (44)
$\geq{L}_{\alpha}\biggl{(}\,\sum_{\mu=1}^{M}\omega_{\mu}\,I_{\mu}\biggr{)}$ (45)
Here, the step (44) follows from (33) and the step (44) follows from convexity of the function $X\mapsto{L}_{\alpha}(X)$. Combining (43) with (45) completes the proof of (40).
For a conical $2$-design from collection of $M$ generalized equiangular lines, the paper [28] proved
$\sum_{\mu=1}^{M}\sum_{j=1}^{N_{\mu}}\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})^{2}=S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+\frac{\sigma-S}{d}\ .$ (46)
Substituting (46) into (33) immediately gives (41). $\blacksquare$
The statement of Proposition 1 poses Tsallis-entropy uncertainty relations for measurements assigned to a generalized equiangular measurement. In particular, the inequality (40) is an entropic uncertainty relation with an average Tsallis $\alpha$-entropy. A distinction from the previous considerations is that averaging in the left-hand side of (40) is taken with generally unequal weights. This is a flip side of possibility to manipulate the coefficients of a generalized equiangular measurement. Another formulation of uncertainty relations uses the Rényi entropies.
Proposition 2
Let ${\mathbb{E}}=\bigl{\{}{\mathcal{E}}_{\mu}\bigr{\}}$ be a set of $M$ generalized symmetric POVMs assigned to a generalized equiangular measurement due to (15). For $\alpha\in[1,2]$, we have
$\sum_{\mu=1}^{M}\omega_{\mu}\,R_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})\geq\frac{1}{1-\alpha}\,\ln\biggl{\{}1+(1-\alpha)L_{\alpha}\biggl{(}\frac{\mathrm{tr}(\boldsymbol{\rho}^{2})\!\;d-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}\gamma_{\mu}^{-1}a_{\mu}\omega_{\mu}\biggr{)}\biggr{\}}\,.$ (47)
Let ${\mathcal{P}}=\bigcup\,{\mathcal{P}}_{\mu}$ be a generalized equiangular measurement whose elements form a conical $2$-design with the parameters (22). For $\alpha\in(0,2]$, it holds that
$R_{\alpha}({\mathcal{P}};\boldsymbol{\rho})\geq\frac{1}{1-\alpha}\,\ln\Bigl{\{}1+(1-\alpha)L_{\alpha}\bigl{(}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+(\sigma-S)d^{-1}\bigr{)}\Bigr{\}}\,.$ (48)
Proof. To prove (47), one writes
$\sum_{\mu=1}^{M}\omega_{\mu}\,R_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})$ $=\frac{1}{1-\alpha}\,\sum_{\mu=1}^{M}\omega_{\mu}\ln\bigl{\{}1+(1-\alpha)H_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})\bigr{\}}$ (49)
$\geq\frac{1}{1-\alpha}\,\sum_{\mu=1}^{M}\omega_{\mu}\ln\bigl{\{}1+(1-\alpha)L_{\alpha}\bigl{(}I_{\mu}\bigr{)}\bigr{\}}\,.$ (50)
Here, the step (49) follows from (28) and the step (50) follows from (33). The formula (50) remains valid for $\alpha\in(0,2]$. If the function $X\mapsto{f}(X)$ is convex, and the function $Y\mapsto{g}(Y)$ is increasing and convex, then the composition $X\mapsto{g}\bigl{(}f(X)\bigr{)}$ is convex as well. Then the features of $X\mapsto{L}_{\alpha}(X)$ and $Y\mapsto(1-\alpha)^{-1}\ln\bigl{\{}1+(1-\alpha)Y\bigr{\}}$ imply convexity of the function [55]
$X\mapsto\frac{1}{1-\alpha}\,\ln\bigl{\{}1+(1-\alpha)L_{\alpha}(X)\bigr{\}}$ (51)
for $\alpha\in[1,2]$. In this interval, the inequality (50) leads to (47).
When the given generalized equiangular measurement is a conical $2$-design, the index of coincidence is fixed by (46) that results in (41). Combining the latter with (28) and increase of the function $Y\mapsto(1-\alpha)^{-1}\ln\bigl{\{}1+(1-\alpha)Y\bigr{\}}$ immediately leads to (48). $\blacksquare$
The inequalities (47) and (48) are Rényi-entropy uncertainty relations for measurements assigned to a generalized equiangular measurement. The first formula evaluates from below the Rényi entropy averaged over POVMs with the actual weights (37). In contrast to (48), it is proved only for $\alpha\in[1,2]$. The latter follows from convexity properties of the functions involved into consideration. The inequality (48) deals with a single generalized equiangular measurement. It holds for $\alpha\in(0,2]$ similarly to (41).
III.2 Fine-grained uncertainty relations
For a pair of observables, uncertainty relations of the Landau–Pollak type are expressed in terms of the maximal probabilities. Its quantum-mechanical interpretation was emphasized in [34], since the original formulation [66] concerned signal analysis. Relations of the Landau–Pollak type can also be considered as fine-grained uncertainty relations. The authors of [67] discussed the role of such relations dealing with a particular combination of the outcomes. This paper deals with POVMs such that the number of outcomes exceeds dimensionality. Then a nontrivial bound from above holds already for a single probability [48]. Namely, the following statement takes place.
Proposition 3
Let ${\mathbb{E}}=\bigl{\{}{\mathcal{E}}_{\mu}\bigr{\}}$ be a set of $M$ generalized symmetric POVMs assigned to a generalized equiangular measurement due to (15), and let $N_{\mu}=N$ for all $\mu=1,\ldots,M$. It holds that
$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{1\leq{j}\leq{N}}{\max}\,\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})\leq\frac{1}{N}\,\biggl{\{}1+\sqrt{N-1}\,\biggl{(}\frac{Nd\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-N}{\mathcal{C}{d}}+\frac{N}{d}\,\sum_{\mu=1}^{M}\gamma_{\mu}^{-1}a_{\mu}\omega_{\mu}-1\biggr{)}^{\!1/2}\,\biggr{\}}\,.$ (52)
Let ${\mathcal{P}}=\bigcup\,{\mathcal{P}}_{\mu}$ be a generalized equiangular measurement whose elements form a conical $2$-design with the parameters (22). The maximal probability satisfies
$\varLambda_{p}\bigl{(}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+(\sigma-S)d^{-1}\bigr{)}\leq\underset{\mu,j}{\max}\,\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})\leq\frac{1}{N_{M}}\,\biggl{\{}1+\sqrt{N_{M}-1}\,\Bigl{(}N_{M}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+N_{M}(\sigma-S)d^{-1}-1\Bigr{)}^{\!1/2}\,\biggr{\}}\,,$ (53)
where the total number of outcomes
$N_{M}=\sum_{\mu=1}^{M}N_{\mu}\,.$ (54)
Proof. The right-hand side of (35) is a concave function of $I(P)$. For a set of $M$ POVMs with the same number of outcomes, we can write
$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{1\leq{j}\leq{N}}{\max}\,\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})\leq\sum_{\mu=1}^{M}\frac{\omega_{\mu}}{N}\,\Bigl{(}1+\sqrt{N-1}\,\sqrt{NI_{\mu}-1}\,\Bigr{)}\leq\frac{1}{N}\,\biggl{\{}1+\sqrt{N-1}\,\biggl{(}N\sum_{\mu=1}^{M}\omega_{\mu}I_{\mu}-1\biggr{)}^{\!1/2}\,\biggr{\}}\,.$ (55)
Combining the latter with (43) completes the proof of (52). The complementarity relation (53) is obtained by substituting (46) into the two-sided estimate (35). $\blacksquare$
The statement of Proposition 3 imposes restrictions on maximal probabilities of the considered measurement. The result (52) estimates from above the maximal probability averaged over POVMs with the weights (37). In general, these weight are not equal. The second result (53) provides estimations of the maximal probability from below and above simultaneously. These results could be converted into inequalities with the corresponding min-entropies. We refrain from presenting the details here.
To formulate relations for the maximal sum of two probabilities, we introduce auxiliary function
$F_{N}(X)=\begin{cases}N^{-1}\bigl{(}2+\sqrt{2N-4}\,\sqrt{NX-1}\,\bigr{)}\,,&\text{if}\ N^{-1}\leq{X}\leq 2^{-1}\,,\\ 1\,,&\text{if}\ 2^{-1}\leq{X}\leq 1\,.\end{cases}$ (56)
One shall now apply (36) to measurements assigned to a generalized equiangular measurement.
Proposition 4
Let ${\mathbb{E}}=\bigl{\{}{\mathcal{E}}_{\mu}\bigr{\}}$ be a set of $M$ generalized symmetric POVMs assigned to a generalized equiangular measurement due to (15), and let $N_{\mu}=N$ for all $\mu=1,\ldots,M$. It holds that
$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{i\neq{j}}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{E}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})\bigr{\}}\leq{F}_{N}\biggl{(}\frac{d\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}\gamma_{\mu}^{-1}a_{\mu}\omega_{\mu}\biggr{)}\,.$ (57)
Let ${\mathcal{P}}=\bigcup\,{\mathcal{P}}_{\mu}$ be a generalized equiangular measurement whose elements form a conical $2$-design with the parameters (22); it holds that
$\underset{(\mu,i)\neq(\nu,j)}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{P}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{P}}_{\nu,j}\boldsymbol{\rho})\bigr{\}}\leq{F}_{N_{M}}\!\Bigl{(}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+(\sigma-S)d^{-1}\Bigr{)}\,,$ (58)
where the total number of outcomes $N_{M}$ is defined by (54).
Proof. By construction, the function $X\mapsto{F}_{N}(X)$ is concave. Due to (36), we therefore have
$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{i\neq{j}}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{E}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})\bigr{\}}\leq\sum_{\mu=1}^{M}\omega_{\mu}\,F_{N}(I_{\mu})\leq{F}_{N}\biggl{(}\,\sum_{\mu=1}^{M}\omega_{\mu}\,I_{\mu}\biggr{)}\,.$ (59)
Combining the latter with (43) completes the proof of (57). The result (58) directly follows from (36) and (46). $\blacksquare$
The statement of Proposition 4 is an example of fine-grained uncertainty relations with the use of probabilities of two separate outcomes. It is not a surprise that the inequality (57) deals with averaging with the weights (37) generally unequal. In the paper [55], relations of such a kind were formulated for measurements assigned to a projective $2$-design. In particular, the case of $d+1$ MUBs was addressed assuming that $d$ is a prime power. Thus, Proposition 4 provides an extension of recent results to generalized equiangular measurements. It would be interesting to examine applications of these inequalities in quantum information, including schemes to detect entanglement and steerability.
III.3 Examples
Let us illustrate the derived relations with some examples. It is instructive to begin with a qubit example. Following [28], we consider the operators
${\mathsf{E}}_{1,1}$ $=\begin{pmatrix}1&0\\ 0&0\end{pmatrix},\qquad{\mathsf{E}}_{1,2}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix},$ (60)
${\mathsf{E}}_{2,1}$ $=\frac{1}{3}\begin{pmatrix}1&-\!\;\mathtt{i}\\ \mathtt{i}&1\end{pmatrix},\qquad{\mathsf{E}}_{2,2}=\frac{1}{6}\begin{pmatrix}2&\sqrt{3}+\mathtt{i}\\ \sqrt{3}-\mathtt{i}&2\end{pmatrix},\qquad{\mathsf{E}}_{2,3}=\frac{1}{6}\begin{pmatrix}2&-\!\;\sqrt{3}+\mathtt{i}\\ -\!\;\sqrt{3}-\mathtt{i}&2\end{pmatrix}.$ (61)
It can be checked that the parameters (10)–(13) read as
$w_{1}=1\,,\qquad w_{2}=\frac{2}{3}\,,\qquad x_{1}=1\,,\qquad x_{2}=\frac{4}{9}\ ,\qquad y_{1}=0\ ,\qquad y_{2}=\frac{1}{9}\ ,\qquad z_{12}=\frac{1}{3}\,.$ (62)
Let us define the five elements of generalized equiangular measurement as
${\mathsf{P}}_{1,i}=\frac{2}{5}\,{\mathsf{E}}_{1,i}\,,\qquad{\mathsf{P}}_{2,j}=\frac{3}{5}\,{\mathsf{E}}_{2,j}\,.$ (63)
We further calculate
$a_{1}=a_{2}=\frac{2}{5}\,,\qquad\mathrm{tr}({\mathsf{P}}_{\mu,j}^{2})=\frac{4}{25}\,,\qquad b_{1}=b_{2}=1\,,\qquad c_{1}=0\,,\qquad c_{2}=\frac{1}{4}\ ,\qquad\omega_{1}=\frac{1}{4}\ ,\qquad\omega_{2}=\frac{3}{4}\ ,$ (64)
and $\mathcal{C}=4$. These values should be kept in mind when looking at the graphs.
Figure 1: Values of the Tsallis $0.8$-entropy versus square of the Bloch vector, whose direction coincides with the $z$-axis on the left and the $x$-axis on the right.
It is instructive to compare the entropies of interest for two POVMs $\bigl{\{}{\mathsf{E}}_{1,1},{\mathsf{E}}_{1,2}\bigr{\}}$ and $\bigl{\{}{\mathsf{E}}_{2,1},{\mathsf{E}}_{2,2},{\mathsf{E}}_{2,3}\bigr{\}}$ separately, the average one and the lower bound (40). These entropic values are shown in Fig. 1 for $\alpha=0.8$ and the two directions of the Bloch vector. The abscissa shows square of the Bloch vector. Of course, the curve of the right-hand side of (40) is the same in both the left and right boxes. It is seen that the lower bound is saturated for the maximally mixed state. The actual direction of the Bloch vector significantly governs the two entropies $H_{\alpha}({\mathcal{E}}_{1};\boldsymbol{\rho})$ and $H_{\alpha}({\mathcal{E}}_{2};\boldsymbol{\rho})$. At the same time, the average entropic value is affected not so essentially. The derived lower bound seems to be sufficiently tight. The deviation is maximal for pure states. On a relative scale, it is around 19.3 % on the left and 14.5 % on the right. The former is particularly conditioned by the fact that $H_{\alpha}({\mathcal{E}}_{1};\boldsymbol{\rho})$ vanishes. With growth of mixedness, the deviation reduces. This example is interesting due to a combination of the two POVMs with different numbers of outcomes. Despite of this choice, the inequality (40) is not far from optimality.
To exemplify the case of conical $2$-design, we take the operators [28]
${\mathsf{P}}_{1,1}=\frac{3-\sqrt{5}}{8}\begin{pmatrix}\sqrt{5}-\sqrt{3}&0\\ 0&\sqrt{5}+\sqrt{3}\end{pmatrix},\qquad{\mathsf{P}}_{1,2}=\frac{3-\sqrt{5}}{8}\begin{pmatrix}\sqrt{5}+\sqrt{3}&0\\ 0&\sqrt{5}-\sqrt{3}\end{pmatrix},$ (65)
${\mathsf{P}}_{2,1}=\frac{3-\sqrt{5}}{8}\begin{pmatrix}2&q\\ q^{*}&2\end{pmatrix},\qquad{\mathsf{P}}_{2,2}=\frac{3-\sqrt{5}}{8}\begin{pmatrix}2&-\!\;\mathtt{i}{q}^{*}\\ \mathtt{i}{q}&2\end{pmatrix},\qquad{\mathsf{P}}_{2,3}=\frac{3-\sqrt{5}}{4\sqrt{2}}\begin{pmatrix}\sqrt{2}&1-\mathtt{i}\\ 1+\mathtt{i}&\sqrt{2}\end{pmatrix}.$ (66)
where $-\!\;q=\bigl{(}2+\sqrt{3}+\mathtt{i}\bigr{)}\sqrt{2-\sqrt{3}}$. The required parameters read as
$\gamma_{1}$ $=\frac{3\sqrt{5}-5}{4}\ ,\qquad\gamma_{1}=1-\gamma_{1}\,,$ (67)
$\mathrm{tr}({\mathsf{P}}_{\mu,j}^{2})$ $=a_{\mu}^{2}b_{\mu}=\frac{7-3\sqrt{5}}{2}\ ,$ (68)
$\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{P}}_{\mu,j})$ $=a_{\mu}^{2}c_{\mu}=\frac{7-3\sqrt{5}}{8}\qquad(i\neq{j})\,,$ (69)
$\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{P}}_{\nu,j})$ $=\frac{7\sqrt{5}-15}{8}\qquad(\mu\neq\nu)\,.$ (70)
It is interesting that the traces (68)–(70) are independent of $\mu$ [28]. Therefore, we obtain here a conical $2$-design such that
$S=\frac{21-9\sqrt{5}}{8}\ ,\qquad\sigma=\frac{11\bigl{(}3-\sqrt{5}\bigr{)}^{2}}{16}\ .$ (71)
Dealing with a single measurement, one can rescale each of the used entropies by its maximal possible value. Then the curves all reduce to the same interval. This example includes five outcomes, so that we divide $H_{\alpha}({\mathcal{P}};\boldsymbol{\rho})$ by $\ln_{\alpha}(5)$ and $R_{\alpha}({\mathcal{P}};\boldsymbol{\rho})$ by $\ln 5$. The rescaled values of the $\alpha$-entropies versus square of the Bloch vector are shown in Fig. 2 for $\alpha=0.8$ and the three directions. The lower bounds are calculated due to (41) and (48), respectively Similarly to the picture of Fig. 1, the deviation from the lower bounds is maximal for pure states. For the Bloch vector along the $z$-axis, it is around 9.1 % in the Tsallis case and 7.9 % in the Rényi case. In a relative scale, deviations are less than for the curves in Fig. 1. Indeed, this example deals with the number of outcomes that is larger.
Figure 2: Rescaled values of the $0.8$-entropies versus square of the Bloch vector for three directions: the Tsallis $0.8$-entropy on the left and the Rényi $0.8$-entropy on the right.
To illustrate some of the results for separate probabilities, we also recall the concept of mutually unbiased measurements (MUMs) [26]. Each of them contains $N_{\mu}=d$ elements. The set =${\mathbb{Q}}=\{{\mathcal{Q}}_{1},\ldots,{\mathcal{Q}}_{M}\}$ is a set of $M$ MUMs of the efficiency $\varkappa$ in dimension $d$, when
$\mathrm{tr}({\mathsf{Q}}_{\mu,j})=1\,,\qquad\mathrm{tr}({\mathsf{Q}}_{\mu,i}{\mathsf{Q}}_{\mu,j})=\delta_{ij}\,\varkappa+(1-\delta_{ij})\>\frac{1-\varkappa}{d-1}\ ,\qquad\mathrm{tr}({\mathsf{Q}}_{\mu,i}{\mathsf{Q}}_{\nu,j})=\frac{1}{d}\qquad(\mu\neq\nu)\,.$ (72)
The authors of [26] showed how to build a complete set of $d+1$ mutually unbiased measurements in ${\mathcal{H}}_{d}$. In general, the parameter $\varkappa$ obeys $d^{-1}\leq\varkappa\leq 1$. The maximal efficiency $\varkappa=1$ corresponds to mutually unbiased bases (MUBs) with rank-one elements. In the considered example, the parameters (10)–(13) read as
$w_{\mu}=1\,,\qquad x_{\mu}=\varkappa\,,\qquad y_{\mu}=\frac{1-\varkappa}{d-1}\ ,\qquad z_{\mu\nu}=\frac{1}{d}\ .$ (73)
It was proved in [50] that
$\frac{1}{M}\,\sum_{\mu=1}^{M}I({\mathcal{Q}}_{\mu};\boldsymbol{\rho})\leq\frac{M-1}{Md}+\frac{1-\varkappa+(\varkappa{d}-1)\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})}{M(d-1)}\ .$ (74)
For $M=d+1$, this inequality is saturated so that [50]
$\frac{1}{d+1}\,\sum_{\mu=1}^{d+1}I({\mathcal{Q}}_{\mu};\boldsymbol{\rho})=\frac{1}{d+1}+\frac{1-\varkappa+(\varkappa{d}-1)\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})}{d^{2}-1}\ .$ (75)
For a set of $d+1$ MUMs, we obviously have $\omega_{\mu}=(d+1)^{-1}$, so that the inequality (52) gives
$\frac{1}{d+1}\,\sum_{\mu=1}^{d+1}\underset{1\leq{j}\leq{d}}{\max}\,\mathrm{tr}({\mathsf{Q}}_{\mu,j}\boldsymbol{\rho})\leq\frac{1}{d}+\frac{1}{d}\,\sqrt{\frac{(\varkappa{d}-1)\bigl{[}d\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-1\bigr{]}}{d+1}}\ .$ (76)
Further, the inequality (57) leads to
$\frac{1}{d+1}\,\sum_{\mu=1}^{d+1}\underset{i\neq{j}}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{Q}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{Q}}_{\mu,j}\boldsymbol{\rho})\bigr{\}}\leq\frac{2}{d}+\frac{\sqrt{2d-4}}{d}\,\sqrt{\frac{(\varkappa{d}-1)\bigl{[}d\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-1\bigr{]}}{d^{2}-1}}\ .$ (77)
For the maximal efficiency $\varkappa=1$ corresponding to MUBs the inequality (77) reduces to one of the results proved in [55]. Then its right-hand side is strictly less than 1 for $d\geq 3$ and an impure state. For $\varkappa<1$ and $d\geq 3$, the inequality (77) is nontrivial for all states. Using MUMs, a generalized equiangular measurement can be built due to (15). We refrain from presenting the details here.
IV Conclusions
We have derived uncertainty relations for measurements assigned to a generalized equiangular measurement. Each of such POVMs is built from generalized equiangular tight frames that are complementary to one another. This concept recently proposed in [28] allows us to manipulate with the involved coefficients in several ways. For instance, there are generalized equiangular measurement for which three out of four defining parameters becomes independent of the measurement operator indices. The role of measurements with a special inner structure is well known in quantum information theory. Complete sets of MUBs and SIC-POVMs are probably most known in this regard. At the same time, the question of existence for such measurements seems to be very hard. To mitigate these questions, one may to weak some of the imposed conditions. Mutually unbiased measurements [26] and general SIC-POVMs [26] are examples of such a kind. The paper [28] proposes further development in this direction.
The presented results follow from an estimation of the index of coincidence for POVMs of interest. In addition, several properties of information characteristics were used to derive uncertainty relations. The used approach is based on inequalities obtained by means of information diagrams. In more detail, these questions are considered in [54, 55]. Namely, generalized entropies of both the Rényi and Tsallis types were utilized. Uncertainty relations with probabilities of separate outcomes have also been addressed. Dealing with average characteristics involves generally unequal weights. The latter is a flip side of existing freedom to manipulate the defining parameters. All the new uncertainty relations are briefly exemplified. Like MUBs and SICs, overcomplete measurements from generalized equiangular tight frames can allow one to detect entanglement and steerability. Applications of the presented uncertainty relations for such purposes will be considered in future investigations.

### v3

@@
-Uncertainty relations for overcomplete measurements from generalized equiangular tight frames
+Uncertainty relations for quantum measurements from generalized equiangular tight frames
 Abstract
-The current study aims to examine uncertainty relations for overcomplete measurements from generalized equiangular tight frames.
+The current study aims to examine uncertainty relations for measurements from generalized equiangular tight frames.
 Informationally overcomplete measurements are a valuable tool in quantum information processing, including tomography and state estimation.
-The maximal sets of mutually unbiased bases are the most common case of informationally overcomplete measurements.
-The existence of $d+1$ of mutually unbiased bases is proved for $d$ being a prime power.
+The maximal sets of mutually unbiased bases are the most common case of such measurements.
+The existence of $d+1$ mutually unbiased bases is proved for $d$ being a prime power.
 More general classes of informationally overcomplete measurements have been proposed for various purposes.
-Such measurements are typically characterized by some inner structure maintaining the required properties.
+Measurements of interest are typically characterized by some inner structure maintaining the required properties.
 It leads to restrictions imposed on generated probabilities.
-To apply the measurements of interest, these restrictions should be converted into information-theoretic terms.
+To apply the considered measurements, these restrictions should be converted into information-theoretic terms.
+It is interesting that certain restrictions hold irrespectively to overcompleteness.
 To describe the amount of uncertainty quantitatively, we use the Tsallis and Rényi entropies as well as probabilities of separate outcomes.
@@
 The derived relations are briefly exemplified.
-overcomplete measurements, index of coincidence, Tsallis entropy, Rényi entropy, conical designs
+symmetric measurements, Tsallis entropy, Rényi entropy, conical designs
 I Introduction
@@
 In fact, the problem of building desired structures in Hilbert spaces is often hard to resolve [8, 16].
-It is known that $d+1$ of mutually unbiased bases in dimension $d$ exist for $d$ being a prime power.
+It is known that $d+1$ mutually unbiased bases in dimension $d$ exist for $d$ being a prime power.
 But the maximal number of mutually unbiased bases remains unknown even for $d=6$ [17].
@@
 In this regard, equiangular tight frames are shown to deserve more attention than they have obtained [19].
-Equiangular measurements are a special case of more general concepts such as equioverlapping measurements [20, 21] and the so-called $(1,M)$-POVMs [22].
+Equiangular measurements are a special case of more general concepts such as equioverlapping measurements [20, 21] and the so-called $(N,M)$-POVMs [22, 23].
 Another way is to allow measurement operators of arbitrary rank.
@@
 The well-known Maassen–Uffink uncertainty relation [35] was originally conjectured by Kraus [36].
-This relation becomes a tool in building feasible schemes to detect of nonclassical correlations such as entanglement [36, 37] and steerability [38, 39, 40].
+This relation becomes a tool in building feasible schemes to detect nonclassical correlations such as entanglement [37, 38] and steerability [39, 40, 41].
 Entropic uncertainty relations have also emerged in the security analysis of quantum cryptographic protocols [42, 43, 44].
@@
 It is important to examine properties of specially built quantum measurements as completely as possible, even if the final opinion on their use is now unformed.
-The aim of this study is to derive uncertainty relations for overcomplete measurements from generalized equiangular tight frames.
-It immediately concerns uncertainty relations for generalized symmetric measurements due to the correspondence mentioned in [28].
-The new relations are expressed in terms of the Renyi [52] and Tsallis entropies [53] as well as sums of probabilities.
+The aim of this study is to derive uncertainty relations for quantum measurements from generalized equiangular tight frames.
+In contrast to the treatment of [29], these measurements are not assumed to be overcomplete.
+The new relations are expressed in terms of the Rényi [53] and Tsallis entropies [54] as well as sums of probabilities.
 Some of the used inequalities are based on considering information diagrams as described in [55, 56].
@@
 Section IV concludes the paper with a summary of the results.
+Appendix A deals with a statement concerning indices of coincidence.
 II Preliminaries
-This section aims to recall the required material concerning informationally overcomplete measurements derived from generalized equiangular tight frames.
+This section aims to recall the required material concerning quantum measurements derived from generalized equiangular tight frames.
 It also considers entropies that will be used to characterize uncertainties in quantum measurements of interest.
@@
 The state of a quantum system is described by density matrix $\boldsymbol{\rho}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ normalized as $\mathrm{tr}(\boldsymbol{\rho})=1$.
-In general, any quantum measurement is represented as a set =${\mathcal{E}}=\bigl{\{}{\mathsf{E}}_{j}\bigr{\}}_{j=1}^{N}$ with ${\mathsf{E}}_{j}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ such that the completeness relation holds,
+In general, any quantum measurement is represented as a set =${\mathcal{E}}=\bigl{\{}{\mathsf{E}}_{j}\bigr{\}}_{j=1}^{N}$ with ${\mathsf{E}}_{j}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ such that the completeness relation holds [57]
 […]
 For the pre-measurement state $\boldsymbol{\rho}$, the probability of $j$-th outcome is equal to $\mathrm{tr}({\mathsf{E}}_{j}\boldsymbol{\rho})$.
-In principle, the number of different outcomes can exceed the dimensionality.
-This property is indispensable in many questions of quantum information science.
+It is very important that the number of different outcomes can exceed the dimensionality.
 A collection of $M$ generalized equiangular lines ${\mathcal{P}}_{\mu}=\bigl{\{}{\mathsf{P}}_{\mu,j}\bigr{\}}_{j=1}^{N_{\mu}}$ with ${\mathsf{P}}_{\mu,j}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ is a generalized equiangular measurement if the following items hold [29]:
@@
 $\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{P}}_{\mu,j})$ $=c_{\mu}\mathrm{tr}({\mathsf{P}}_{\mu,i})\,\mathrm{tr}({\mathsf{P}}_{\mu,j})\qquad(i\neq{j})\,,$ (6)
-$\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{P}}_{\nu,j})$ $=f\,\mathrm{tr}({\mathsf{P}}_{\mu,j})\,\mathrm{tr}({\mathsf{P}}_{\nu,j})\qquad(\mu\neq\nu)\,.$ (7)
+$\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{P}}_{\nu,j})$ $=f\,\mathrm{tr}({\mathsf{P}}_{\mu,i})\,\mathrm{tr}({\mathsf{P}}_{\nu,j})\qquad(\mu\neq\nu)\,.$ (7)
 From the above definition, it follows that $f=d^{-1}$,
 $a_{\mu}=\frac{\gamma_{\mu}d}{N_{\mu}}\ ,\qquad c_{\mu}=\frac{N_{\mu}f-b_{\mu}}{N_{\mu}-1}\ .$ (8)
-The admissible range for $b_{\mu}$ is obtained as [28]
-$\frac{1}{d}<b_{\mu}\leq\frac{\min\{d,N_{\mu}\}}{d}\ .$ (9)
-A single equiangular measurement is a particular case for $M=\gamma=b=1$.
-It should be noted here that our notation slightly differs from the notation of [28].
+A single equiangular measurement with rank-one elements is a particular case for $M=\gamma=b=1$.
 The generalized equiangular measurement can also be interpreted as a set of $M$ POVMs so that $\mu$-th POVM has elements $\gamma_{\mu}^{-1}\,{\mathsf{P}}_{\mu,j}$.
@@
 The generalized symmetric measurements were introduced as collections of POVMs that are not equinumerous [28].
-It is important here that elements of generalized equiangular measurements can be constructed from elements of generalized symmetric measurements.
+A generalized equiangular measurement can be constructed from elements of generalized symmetric measurements.
 One uses [28] a set of $M$ POVMs ${\mathcal{E}}_{\mu}=\bigl{\{}{\mathsf{E}}_{\mu,j}\bigr{\}}_{j=1}^{N_{\mu}}$ satisfying additional symmetry conditions:
@@
 $\sum_{\mu=1}^{M}N_{\mu}=d^{2}+M-1\,.$ (13)
-There are four construction methods with the use of linear combinations of Hermitian orthonormal operator bases [27]. We refrain from presenting the details here.
-The following observation has been made in [28].
-Suppose that positive numbers $\gamma_{\mu}$ obey (3) and $M$ POVMs ${\mathcal{E}}_{\mu}=\bigl{\{}{\mathsf{E}}_{\mu,j}\bigr{\}}_{j=1}^{N_{\mu}}$ satisfy (10)–(13).
-Then operators of the form
+Let positive numbers $\gamma_{\mu}$ obey (3) and $M$ POVMs ${\mathcal{E}}_{\mu}=\bigl{\{}{\mathsf{E}}_{\mu,j}\bigr{\}}_{j=1}^{N_{\mu}}$ satisfy (9)–(12). Then operators of the form
 ${\mathsf{P}}_{\mu,j}=\gamma_{\mu}\,{\mathsf{E}}_{\mu,j}$ (14)
-compose a generalized equiangular measurement. The correspondence between the coefficients reads as
+compose a generalized equiangular measurement [29]. The correspondence between the coefficients reads as
 $a_{\mu}=\gamma_{\mu}w_{\mu}\,,\qquad b_{\mu}=\frac{x_{\mu}}{w_{\mu}^{2}}\ ,\qquad c_{\mu}=\frac{y_{\mu}}{w_{\mu}^{2}}\ ,\qquad f=\frac{z_{\mu\nu}}{w_{\mu}w_{\nu}}\ .$ (15)
 The obtained POVM is informationally complete, if it is constructed from informationally complete set of elements ${\mathsf{E}}_{\mu,j}$ and $\gamma_{\mu}>0$ for all $\mu=1,\ldots,M$ [29].
-Thus, the generalized symmetric measurements are sufficient to build a generalized equiangular measurement.
 Despite of simplicity of the relation (14), this construction has some nice features.
-For example, it is possible to manipulate the coefficients in such a way that both the traces $\mathrm{tr}({\mathsf{P}}_{\mu,j}^{2})$ and $\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{P}}_{\mu,j})$ become $\mu$-independent [28].
-Projective designs have found use in various questions of quantum information theory.
-It seems that the special case of $2$-designs is provided with a large number of examples.
-Let $\bigl{\{}|\phi_{j}\rangle\bigr{\}}_{j=1}^{N}$ be a set of $N$ unit vectors of ${\mathcal{H}}_{d}$.
+For example, one can manipulate the coefficients so that both the traces $\mathrm{tr}({\mathsf{P}}_{\mu,j}^{2})$ and $\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{P}}_{\mu,j})$ become $\mu$-independent [29].
+Under additional conditions on the coefficients, a generalized equiangular measurement is a conical $2$-design [29].
+Projective and conical designs have found use in various questions of quantum information theory.
+Let $\bigl{\{}|\phi_{j}\rangle\bigr{\}}_{j=1}^{K}$ be a set of $K$ unit vectors of ${\mathcal{H}}_{d}$.
 There are several equivalent definitions, when these vectors form a projective $2$-design [18, 58].
 One of them reads as
-$\frac{1}{N}\,\sum_{j=1}^{N}|\phi_{j}\rangle\langle\phi_{j}|\otimes|\phi_{j}\rangle\langle\phi_{j}|=\frac{2}{d^{2}+d}\>{\mathsf{\Pi}}_{\mathrm{sym}}\,,$ (17)
+$\frac{1}{K}\,\sum_{j=1}^{K}|\phi_{j}\rangle\langle\phi_{j}|\otimes|\phi_{j}\rangle\langle\phi_{j}|=\frac{2}{d^{2}+d}\>{\mathsf{\Pi}}_{\mathrm{sym}}\,,$ (16)
 where ${\mathsf{\Pi}}_{\mathrm{sym}}$ is the orthogonal projection on the symmetric subspace of ${\mathcal{H}}_{d}\otimes{\mathcal{H}}_{d}$.
@@
 We introduce the unitary swap operator $\tilde{\mathsf{W}}$ on ${\mathcal{H}}_{d}\otimes{\mathcal{H}}_{d}$ which takes $|\psi\rangle\otimes|\varphi\rangle$ to $|\varphi\rangle\otimes|\psi\rangle$.
-The set $\bigl{\{}{\mathsf{A}}_{j}\bigr{\}}_{j=1}^{N}$ with ${\mathsf{A}}_{j}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ is a conical $2$-design, if and only if
+The set $\bigl{\{}{\mathsf{A}}_{j}\bigr{\}}_{j=1}^{K}$ with ${\mathsf{A}}_{j}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ is a conical $2$-design, if and only if
 […]
-where $\kappa_{+}\geq\kappa_{-}>0$. It can be noticed that $2\kappa_{\pm}=\kappa_{\mathrm{s}}\pm\kappa_{\mathrm{a}}$. Taking the partial trace in (18) yields
-[…]
-To each conical $2$-design, we can assign POVM =${\mathcal{E}}=\bigl{\{}{\mathsf{E}}_{j}\bigr{\}}_{j=1}^{N}$ with elements
-${\mathsf{E}}_{j}=\frac{\mathrm{tr}({\mathsf{A}}_{j})}{\kappa_{+}d+\kappa_{-}}\,{\mathsf{A}}_{j}\,.$ (20)
-The case with constant $\mathrm{tr}({\mathsf{A}}_{j})$ is very close to a projective $2$-design. On the other hand, a possibility to manipulate the coefficients generally excludes this restriction.
-The author of [28] showed that some of the generalized equiangular measurements are conical $2$-designs. Let us take the coefficients such that
+where $\kappa_{+}\geq\kappa_{-}>0$.
+It can be noticed that $2\kappa_{\pm}=\kappa_{\mathrm{s}}\pm\kappa_{\mathrm{a}}$.
+To obtain a conical $2$-design from equiangular lines, the coefficients should satisfy [29]
 $a_{\mu}^{2}(b_{\mu}-c_{\mu})=S$ (18)
@@
 There exist conical $2$-designs even though they are constructed from the generalized symmetric measurements that are not conical $2$-designs [29].
-This property could be useful in producing new conical $2$-designs.
-Unitary designs are now considered as a powerful tool in quantum information science [59, 60].
-Thus, generalized equiangular measurements deserve to be studied in application to questions of quantum information science.
+This property is useful in producing new conical $2$-designs.
+Unitary designs are also considered as a powerful tool in quantum information science [61, 62].
 It turned out that generated probabilities satisfy certain conditions that lead to various restrictions including uncertainty relations.
@@
 Let $P=(p_{1},\ldots,p_{N})$ be a probability distribution with the index running over the set $\{1,\ldots,N\}$.
-The corresponding index of coincidence reads as
+The index of coincidence reads as
 $I(P)=\sum_{j=1}^{N}p_{j}^{2}\,.$ (20)
-Indicies with larger degrees of probabilities have also found use [61].
-Applications of such indicies to derive uncertainty relations were given in [62, 63, 64].
+Indices with larger degrees of probabilities have also found use [63].
+Applications of such indices to derive uncertainty relations were given in [64, 65, 66].
 For $\alpha>0$, the Tsallis $\alpha$-entropy is expressed as [54]
@@
 It was demonstrated for information diagrams with the Shannon entropy that the theoretically best lower-bounding curve is not smooth [63].
-Expectedly, the same feature takes place for the Tsallis $\alpha$-entropy.
+Naturally, the same feature takes place for the Tsallis $\alpha$-entropy.
 It can be shown that a smoothness is violated in the points with abscissas $X=k^{-1}$.
@@
 For $\alpha=1$, the inequality (30) reduces to one of the main results of [63].
-We will apply (33) to derive uncertainty relations for informationally overcomplete measurements and conical $2$-designs.
+We will apply (30) to derive uncertainty relations for quantum measurements of interest including conical $2$-designs.
 Information diagrams have also been applied to the maximal probability.
@@
 Each of these points of the diagram corresponds to a uniform probability distribution [55, 56].
-One of the results of [55] allows one to estimate sums of two probabilities from above at the given index of coincidence.
+One of the results of [56] allows us to estimate sums of two probabilities from above at the given index of coincidence.
 Let the index of coincidence of the distribution $P=(p_{1},\ldots,p_{N})$ obey $N^{-1}\leq{I}(P)\leq 2^{-1}$.
@@
 The inequality (33) is formally correct for all acceptable values of $I(P)$, but a nontrivial bound is obtained only for $I(P)<2^{-1}$.
-As a rule, the number of outcomes in overcomplete measurements is sufficiently large.
-Hence, the last inequality is typically satisfied.
-For example, this holds for a complete set of $d+1$ MUBs in dimension $d\geq 3$ [55].
+This inequality takes place for sufficiently large number of outcomes.
+For example, it holds for a complete set of $d+1$ MUBs in dimension $d\geq 3$ [56].
 III Main results
-This section presents uncertainty relations for informationally overcomplete measurements generalized equiangular tight frames including conical $2$-designs.
-One shall address the entropic formulation as well as the case of probabilities for separate outcomes.
+This section presents uncertainty relations for quantum measurements from generalized equiangular tight frames including conical $2$-designs.
+One shall address the entropic formulation as well as the case of probabilities of separate outcomes.
 All the derived relations will be exemplified.
@@
 Statistics of the measurements of interest can be interpreted in two ways.
-First, it relates to a single generalized equiangular measurement.
+First, it deals with a single generalized equiangular measurement.
 Second, one links to a set of generalized symmetric measurements.
@@
 This fact leads to a lot of consequences.
-To avoid bulky expressions, we will use positive numbers
-$\omega_{\mu}$ $=\frac{1}{\mathcal{C}}\>\frac{\gamma_{\mu}^{2}}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}\ ,$ (37)
-$\mathcal{C}$ $=\sum_{\mu=1}^{M}\frac{\gamma_{\mu}^{2}}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}\ .$ (38)
+To avoid bulky expressions, we will use the positive numbers
+$\omega_{\mu}$ $=\frac{\mathcal{C}^{-1}}{x_{\mu}-y_{\mu}}=\frac{\mathcal{C}^{-1}\gamma_{\mu}^{2}}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}\ ,$ (34)
+$\mathcal{C}$ $=\sum_{\mu=1}^{M}\frac{1}{x_{\mu}-y_{\mu}}\ .$ (35)
 By construction, one has
@@
 Let ${\mathbb{E}}=\bigl{\{}{\mathcal{E}}_{\mu}\bigr{\}}$ be a set of $M$ generalized symmetric POVMs assigned to a generalized equiangular measurement due to (14). For $\alpha\in(0,2]$, we have
-$\sum_{\mu=1}^{M}\omega_{\mu}\,H_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})\geq{L}_{\alpha}\biggl{(}\frac{d\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}\gamma_{\mu}^{-1}a_{\mu}\omega_{\mu}\biggr{)}\,.$ (40)
+$\sum_{\mu=1}^{M}\omega_{\mu}\,H_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})\geq{L}_{\alpha}\biggl{(}\frac{d\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}\biggr{)}\,.$ (37)
 Let ${\mathcal{P}}=\bigcup\,{\mathcal{P}}_{\mu}$ be a generalized equiangular measurement whose elements form a conical $2$-design with the parameters (19). For $\alpha\in(0,2]$, it holds that
 $H_{\alpha}({\mathcal{P}};\boldsymbol{\rho})\geq{L}_{\alpha}\bigl{(}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+(\sigma-S)d^{-1}\bigr{)}\,.$ (38)
-Proof. It was shown in appendix C of [28] that
-$\mathrm{tr}(\boldsymbol{\rho}^{2})-\frac{1}{d}=\sum_{\mu=1}^{M}\frac{1}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}\biggl{(}\sum_{j=1}^{N_{\mu}}\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})^{2}-\frac{a_{\mu}\gamma_{\mu}}{d}\biggr{)}.$ (42)
-Combining this with (15) and (37) immediately gives
-$\mathcal{C}\,\sum_{\mu=1}^{M}\omega_{\mu}\,I_{\mu}=\mathrm{tr}(\boldsymbol{\rho}^{2})-\frac{1}{d}+\frac{1}{d}\,\sum_{\mu=1}^{M}\frac{\gamma_{\mu}}{a_{\mu}(b_{\mu}-c_{\mu})}\ ,$ (43)
-where $I_{\mu}\equiv{I}({\mathcal{E}}_{\mu};\boldsymbol{\rho})$. Using (39), we further write
+Proof. To prove (37), we denote (84) by $I_{\mu}$ and further write
 $\sum_{\mu=1}^{M}\omega_{\mu}\,H_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})$ $\geq\sum_{\mu=1}^{M}\omega_{\mu}\,L_{\alpha}(I_{\mu})$ (39)
-$\geq{L}_{\alpha}\biggl{(}\,\sum_{\mu=1}^{M}\omega_{\mu}\,I_{\mu}\biggr{)}$ (45)
-Here, the step (44) follows from (33) and the step (44) follows from convexity of the function $X\mapsto{L}_{\alpha}(X)$. Combining (43) with (45) completes the proof of (40).
-For a conical $2$-design from collection of $M$ generalized equiangular lines, the paper [28] proved
+$\geq{L}_{\alpha}\biggl{(}\,\sum_{\mu=1}^{M}\omega_{\mu}\,I_{\mu}\biggr{)}\,.$ (40)
+Here, the step (39) follows from (30) and the step (40) follows from convexity of the function $X\mapsto{L}_{\alpha}(X)$.
+Combining (85) with (40) completes the proof of (37), since the function $X\mapsto{L}_{\alpha}(X)$ decreases.
+For a conical $2$-design from collection of $M$ generalized equiangular lines, the author of [29] proved that
 $\sum_{\mu=1}^{M}\sum_{j=1}^{N_{\mu}}\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})^{2}=S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+\frac{\sigma-S}{d}\ .$ (41)
 Substituting (41) into (30) immediately gives (38). $\blacksquare$
-The statement of Proposition 1 poses Tsallis-entropy uncertainty relations for measurements assigned to a generalized equiangular measurement.
-In particular, the inequality (40) is an entropic uncertainty relation with an average Tsallis $\alpha$-entropy.
+The statement of Proposition 1 gives Tsallis-entropy uncertainty relations for measurements assigned to a generalized equiangular measurement.
+In particular, the inequality (37) is an entropic uncertainty relation with average Tsallis $\alpha$-entropy.
 A distinction from the previous considerations is that averaging in the left-hand side of (37) is taken with generally unequal weights.
 This is a flip side of possibility to manipulate the coefficients of a generalized equiangular measurement.
-Another formulation of uncertainty relations uses the Rényi entropies.
+In contrast to (37), the inequality (38) deals with a single measurement.
+Instead of (85), the corresponding index of coincidence is calculated exactly due to (41).
+The weights $\omega_{\mu}$ in the left-hand side of (37) are determined by the coefficients of a generalized equiangular measurement.
+Since $\mathrm{tr}(\boldsymbol{\rho}^{2})=1$ for a pure state, the inequalities (37) and (38) then give
+$\sum_{\mu=1}^{M}\omega_{\mu}\,H_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})$ $\geq{L}_{\alpha}\biggl{(}\frac{d-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}\biggr{)}\,,$ (42)
+$H_{\alpha}({\mathcal{P}};\boldsymbol{\rho})$ $\geq{L}_{\alpha}\bigl{(}S+(\sigma-S)d^{-1}\bigr{)}\,.$ (43)
+These relations actually hold for all states, because the function $X\mapsto{L}_{\alpha}(X)$ decreases.
+Such inequalities are typically used in criteria to detect nonclassical correlations.
+Together with the weights $\omega_{\mu}$, the right-hand side of (42) can be treated as an estimation of Tsallis entropies related to $M$ generalized symmetric POVMs.
+In a similar vein, the right-hand side of (43) is an entropic characteristic of the given conical $2$-design.
+The next formulation of uncertainty relations uses the Rényi entropies.
 Proposition 2
 Let ${\mathbb{E}}=\bigl{\{}{\mathcal{E}}_{\mu}\bigr{\}}$ be a set of $M$ generalized symmetric POVMs assigned to a generalized equiangular measurement due to (14). For $\alpha\in[1,2]$, we have
-$\sum_{\mu=1}^{M}\omega_{\mu}\,R_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})\geq\frac{1}{1-\alpha}\,\ln\biggl{\{}1+(1-\alpha)L_{\alpha}\biggl{(}\frac{\mathrm{tr}(\boldsymbol{\rho}^{2})\!\;d-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}\gamma_{\mu}^{-1}a_{\mu}\omega_{\mu}\biggr{)}\biggr{\}}\,.$ (47)
+$\sum_{\mu=1}^{M}\omega_{\mu}\,R_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})\geq\frac{1}{1-\alpha}\,\ln\biggl{\{}1+(1-\alpha)L_{\alpha}\biggl{(}\frac{\mathrm{tr}(\boldsymbol{\rho}^{2})\!\;d-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}\biggr{)}\biggr{\}}\,.$ (44)
 Let ${\mathcal{P}}=\bigcup\,{\mathcal{P}}_{\mu}$ be a generalized equiangular measurement whose elements form a conical $2$-design with the parameters (19). For $\alpha\in(0,2]$, it holds that
@@
 $X\mapsto\frac{1}{1-\alpha}\,\ln\bigl{\{}1+(1-\alpha)L_{\alpha}(X)\bigr{\}}$ (48)
-for $\alpha\in[1,2]$. In this interval, the inequality (50) leads to (47).
+for $\alpha\in[1,2]$. In this range, the inequalities (47) and (85) lead to (44), since the function (48) decreases.
 When the given generalized equiangular measurement is a conical $2$-design, the index of coincidence is fixed by (41) that results in (38).
@@
 The inequalities (44) and (45) are Rényi-entropy uncertainty relations for measurements assigned to a generalized equiangular measurement.
-The first formula evaluates from below the Rényi entropy averaged over POVMs with the actual weights (37).
+The first formula estimates from below the Rényi entropy averaged over POVMs with the actual weights (34).
 In contrast to (45), it is proved only for $\alpha\in[1,2]$.
@@
 It holds for $\alpha\in(0,2]$ similarly to (38).
+For a pure state, the inequalities (44) and (45) give
+$\sum_{\mu=1}^{M}\omega_{\mu}\,R_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})$ $\geq\frac{1}{1-\alpha}\,\ln\biggl{\{}1+(1-\alpha)L_{\alpha}\biggl{(}\frac{d-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}\biggr{)}\biggr{\}}\,,$ (49)
+$R_{\alpha}({\mathcal{P}};\boldsymbol{\rho})$ $\geq\frac{1}{1-\alpha}\,\ln\Bigl{\{}1+(1-\alpha)L_{\alpha}\bigl{(}S+(\sigma-S)d^{-1}\bigr{)}\Bigr{\}}\,.$ (50)
+Similarly to (42) and (43), these relations hold for all states and have potential applications.
+The right-hand sides of (49) and (50) can be interpreted as an estimate of Rényi entropies assigned to the corresponding measurements.
 III.2 Fine-grained uncertainty relations
@@
 It holds that
-$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{1\leq{j}\leq{N}}{\max}\,\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})\leq\frac{1}{N}\,\biggl{\{}1+\sqrt{N-1}\,\biggl{(}\frac{Nd\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-N}{\mathcal{C}{d}}+\frac{N}{d}\,\sum_{\mu=1}^{M}\gamma_{\mu}^{-1}a_{\mu}\omega_{\mu}-1\biggr{)}^{\!1/2}\,\biggr{\}}\,.$ (52)
+$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{1\leq{j}\leq{N}}{\max}\,\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})\leq\frac{1}{N}\,\biggl{\{}1+\sqrt{N-1}\,\biggl{(}\frac{Nd\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-N}{\mathcal{C}{d}}+\frac{N}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}-1\biggr{)}^{\!1/2}\,\biggr{\}}\,.$ (51)
 Let ${\mathcal{P}}=\bigcup\,{\mathcal{P}}_{\mu}$ be a generalized equiangular measurement whose elements form a conical $2$-design with the parameters (19). The maximal probability satisfies
-$\varLambda_{p}\bigl{(}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+(\sigma-S)d^{-1}\bigr{)}\leq\underset{\mu,j}{\max}\,\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})\leq\frac{1}{N_{M}}\,\biggl{\{}1+\sqrt{N_{M}-1}\,\Bigl{(}N_{M}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+N_{M}(\sigma-S)d^{-1}-1\Bigr{)}^{\!1/2}\,\biggr{\}}\,,$ (53)
+$\varLambda_{p}\bigl{(}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+(\sigma-S)d^{-1}\bigr{)}\leq\underset{\mu,j}{\max}\,\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})\leq\frac{1}{K}\,\biggl{\{}1+\sqrt{K-1}\,\Bigl{(}KS\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+K(\sigma-S)d^{-1}-1\Bigr{)}^{\!1/2}\,\biggr{\}}\,,$ (52)
 where the total number of outcomes
-$N_{M}=\sum_{\mu=1}^{M}N_{\mu}\,.$ (54)
+$K=\sum_{\mu=1}^{M}N_{\mu}\,.$ (53)
 Proof. The right-hand side of (32) is a concave function of $I(P)$. For a set of $M$ POVMs with the same number of outcomes, we can write
@@
 We refrain from presenting the details here.
+The state-independent counterparts of (51) and (52) respectively read as
+$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{1\leq{j}\leq{N}}{\max}\,\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})$ $\leq\frac{1}{N}\,\biggl{\{}1+\sqrt{N-1}\,\biggl{(}\frac{Nd-N}{\mathcal{C}{d}}+\frac{N}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}-1\biggr{)}^{\!1/2}\,\biggr{\}}\,,$ (55)
+$\underset{\mu,j}{\max}\,\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})$ $\leq\frac{1}{K}\,\biggl{\{}1+\sqrt{K-1}\,\Bigl{(}KS+K(\sigma-S)d^{-1}-1\Bigr{)}^{\!1/2}\,\biggr{\}}\,.$ (56)
+These inequalities are obtained by replacing $\mathrm{tr}(\boldsymbol{\rho}^{2})$ with its maximal value one.
+Note that only an estimate from above appears in (56).
+The formulas (55) and (56) allow us to estimate the maximal probabilities from above in terms of the characteristics of POVMs solely.
 To formulate relations for the maximal sum of two probabilities, we introduce auxiliary function
@@
 It holds that
-$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{i\neq{j}}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{E}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})\bigr{\}}\leq{F}_{N}\biggl{(}\frac{d\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}\gamma_{\mu}^{-1}a_{\mu}\omega_{\mu}\biggr{)}\,.$ (57)
+$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{i\neq{j}}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{E}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})\bigr{\}}\leq{F}_{N}\biggl{(}\frac{d\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}\biggr{)}\,.$ (58)
 Let ${\mathcal{P}}=\bigcup\,{\mathcal{P}}_{\mu}$ be a generalized equiangular measurement whose elements form a conical $2$-design with the parameters (19); it holds that
-$\underset{(\mu,i)\neq(\nu,j)}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{P}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{P}}_{\nu,j}\boldsymbol{\rho})\bigr{\}}\leq{F}_{N_{M}}\!\Bigl{(}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+(\sigma-S)d^{-1}\Bigr{)}\,,$ (58)
-where the total number of outcomes $N_{M}$ is defined by (54).
-Proof. By construction, the function $X\mapsto{F}_{N}(X)$ is concave. Due to (36), we therefore have
+$\underset{(\mu,i)\neq(\nu,j)}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{P}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{P}}_{\nu,j}\boldsymbol{\rho})\bigr{\}}\leq{F}_{K}\!\bigl{(}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+(\sigma-S)d^{-1}\bigr{)}\,,$ (59)
+where the total number of outcomes $K$ is defined by (53).
+Proof. By construction, the function $X\mapsto{F}_{N}(X)$ is increasing and concave. Due to (33), we therefore have
 $\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{i\neq{j}}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{E}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})\bigr{\}}\leq\sum_{\mu=1}^{M}\omega_{\mu}\,F_{N}(I_{\mu})\leq{F}_{N}\biggl{(}\,\sum_{\mu=1}^{M}\omega_{\mu}\,I_{\mu}\biggr{)}\,.$ (60)
@@
 The statement of Proposition 4 is an example of fine-grained uncertainty relations with the use of probabilities of two separate outcomes.
-It is not a surprise that the inequality (57) deals with averaging with the weights (37) generally unequal.
-In the paper [55], relations of such a kind were formulated for measurements assigned to a projective $2$-design.
+It is not a surprise that the inequality (58) realizes averaging with the weights (34) generally unequal.
+In the paper [56], similar relations with equal weights were formulated for measurements assigned to a projective $2$-design.
 In particular, the case of $d+1$ MUBs was addressed assuming that $d$ is a prime power.
 Thus, Proposition 4 provides an extension of recent results to generalized equiangular measurements.
+To convert (58) and (59) into the state-independent form, we replace $\mathrm{tr}(\boldsymbol{\rho}^{2})$ with its maximal value one, whence
+$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{i\neq{j}}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{E}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})\bigr{\}}\leq{F}_{N}\biggl{(}\frac{d-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}\biggr{)}\,,$ (61)
+$\underset{(\mu,i)\neq(\nu,j)}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{P}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{P}}_{\nu,j}\boldsymbol{\rho})\bigr{\}}\leq{F}_{K}\!\bigl{(}S+(\sigma-S)d^{-1}\bigr{)}\,.$ (62)
 It would be interesting to examine applications of these inequalities in quantum information, including schemes to detect entanglement and steerability.
 III.3 Examples
-Let us illustrate the derived relations with some examples. It is instructive to begin with a qubit example. Following [28], we consider the operators
+Let us illustrate the derived relations with some examples. It is instructive to begin with a qubit. Its density matrix reads as
+[…]
+where $r_{x}=\sin\theta\cos\varphi$, $r_{y}=\sin\theta\sin\varphi$ and $r_{z}=\cos\theta$ are the components of the Bloch vector.
+By $\theta$ and $\varphi$, we mean usual angles on the Bloch sphere.
+Three eigenbases of the Pauli matrices $\boldsymbol{\tau}_{x}$, $\boldsymbol{\tau}_{y}$ and $\boldsymbol{\tau}_{z}$ are mutually unbiased.
+This case deals with three von Neumann measurements.
+Here, the weights (34) are equal to one third.
+The six states of the three eigenbases also form a $3$-design [64] and, herewith, a $2$-design.
+The resulting POVM consists of elements of the form $3^{-1}|b_{\mu,j}\rangle\langle{b}_{\mu,j}|$ with $\mu=x,y,z$ and $j=\pm$.
+It follows from (37) and (38) that, for $\alpha\in[0,2]$,
+$\frac{1}{3}\sum_{\mu=x,y,z}H_{\alpha}({\mathcal{B}}_{\mu};\boldsymbol{\rho})\geq{L}_{\alpha}\biggl{(}\frac{3+r^{2}}{6}\biggr{)}\,,$ (64)
+$H_{\alpha}({\mathcal{P}};\boldsymbol{\rho})\geq{L}_{\alpha}\biggl{(}\frac{3+r^{2}}{18}\biggr{)}\,,$ (65)
+where $r^{2}=r_{x}^{2}+r_{y}^{2}+r_{z}^{2}$ is square of the Bloch vector.
+The uncertainty relations (64) and (65) were actually reported in [56].
+A utility of these inequalities was also discussed therein.
+The next example deals with unequal weights (34). Following [29], we consider the operators
 ${\mathsf{E}}_{1,1}$ $=\begin{pmatrix}1&0\\ 0&0\end{pmatrix},\qquad{\mathsf{E}}_{1,2}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix},$ (66)
@@
 Despite of this choice, the inequality (37) is not far from optimality.
-To exemplify the case of conical $2$-design, we take the operators [28]
+To exemplify the case of conical $2$-design, we use the operators [29]
 ${\mathsf{P}}_{1,1}=\frac{3-\sqrt{5}}{8}\begin{pmatrix}\sqrt{5}-\sqrt{3}&0\\ 0&\sqrt{5}+\sqrt{3}\end{pmatrix},\qquad{\mathsf{P}}_{1,2}=\frac{3-\sqrt{5}}{8}\begin{pmatrix}\sqrt{5}+\sqrt{3}&0\\ 0&\sqrt{5}-\sqrt{3}\end{pmatrix},$ (71)
@@
 where $-\!\;q=\bigl{(}2+\sqrt{3}+\mathtt{i}\bigr{)}\sqrt{2-\sqrt{3}}$. The required parameters read as
-$\gamma_{1}$ $=\frac{3\sqrt{5}-5}{4}\ ,\qquad\gamma_{1}=1-\gamma_{1}\,,$ (67)
+$\gamma_{1}$ $=\frac{3\sqrt{5}-5}{4}\ ,\qquad\gamma_{2}=1-\gamma_{1}\,,$ (73)
 $\mathrm{tr}({\mathsf{P}}_{\mu,j}^{2})$ $=a_{\mu}^{2}b_{\mu}=\frac{7-3\sqrt{5}}{2}\ ,$ (74)
@@
 2 for $\alpha=0.8$ and the three directions.
-The lower bounds are calculated due to (41) and (48), respectively Similarly to the picture of Fig.
+The lower bounds are calculated due to (38) and (45), respectively.
+Similarly to the picture of Fig.
 1, the deviation from the lower bounds is maximal for pure states.
@@
 In general, the parameter $\varkappa$ obeys $d^{-1}\leq\varkappa\leq 1$.
-The maximal efficiency $\varkappa=1$ corresponds to mutually unbiased bases (MUBs) with rank-one elements.
+The maximal efficiency $\varkappa=1$ corresponds to mutually unbiased bases (MUBs).
 In the considered example, the parameters (9)–(12) read as
@@
 For the maximal efficiency $\varkappa=1$ corresponding to MUBs the inequality (83) reduces to one of the results proved in [56].
-Then its right-hand side is strictly less than 1 for $d\geq 3$ and an impure state.
+Then its right-hand side is strictly less than 1 for $d\geq 3$ and a state that is not pure.
 For $\varkappa<1$ and $d\geq 3$, the inequality (83) is nontrivial for all states.
@@
 At the same time, the question of existence for such measurements seems to be very hard.
-To mitigate these questions, one may to weak some of the imposed conditions.
+To mitigate these questions, one can weak some of the imposed conditions.
 Mutually unbiased measurements [27] and general SIC-POVMs [27] are examples of such a kind.
@@
 All the new uncertainty relations are briefly exemplified.
-Like MUBs and SICs, overcomplete measurements from generalized equiangular tight frames can allow one to detect entanglement and steerability.
+Like MUBs and SICs, quantum measurements from generalized equiangular tight frames can allow one to detect entanglement and steerability.
 Applications of the presented uncertainty relations for such purposes will be considered in future investigations.
+Appendix A A statement on indices of coincidence
+In principle, the required technique was developed by the author of [29].
+The key distinction of this appendix is that the considered sets of operators are assumed to be incomplete in general.
+Instead of the equality presented in [29], the following inequality takes place.
+Lemma 1
+Let ${\mathbb{E}}=\bigl{\{}{\mathcal{E}}_{\mu}\bigr{\}}$ be a set of $M$ generalized symmetric POVMs assigned to a generalized equiangular measurement due to (14), with the $\mu$-th index of coincidence
+$I({\mathcal{E}}_{\mu};\boldsymbol{\rho})=\sum_{j=1}^{N_{\mu}}\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})^{2}\,.$ (84)
+It holds that
+$\sum_{\mu=1}^{M}\omega_{\mu}\,I({\mathcal{E}}_{\mu};\boldsymbol{\rho})\leq\frac{d\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}\,,$ (85)
+where the numbers $\omega_{\mu}$ and $\mathcal{C}$ are defined by (34) and (35), respectively.
+Proof. Instead of (85), we aim to prove the equivalent inequality
+$\sum_{\mu=1}^{M}\frac{1}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}\,\biggl{(}\,\sum_{j=1}^{N_{\mu}}\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})^{2}-\frac{a_{\mu}\gamma_{\mu}}{d}\biggr{)}\leq\mathrm{tr}(\boldsymbol{\rho}^{2})-f\ .$ (86)
+Their equivalence can be checked immediately due to (14) and (15). Let us represent an arbitrary $\boldsymbol{\rho}$ as
+$\boldsymbol{\rho}=\sum_{\mu=1}^{M}\sum_{j=1}^{N_{\mu}}\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho}){\mathsf{G}}_{\mu,j}+\boldsymbol{\varpi}\,.$ (87)
+For the given elements ${\mathsf{P}}_{\mu,j}$, the operators ${\mathsf{G}}_{\mu,j}$ read as [29]
+[…]
+By $\boldsymbol{\varpi}$, we mean in (87) potentially nonzero part of $\boldsymbol{\rho}$ such that $\mathrm{tr}({\mathsf{G}}_{\mu,j}\boldsymbol{\varpi})=0$ for all $\mu$ and $j$.
+Being depending on $\boldsymbol{\rho}$, this term does not appear in (86).
+To check (87), one writes
+$\mathrm{tr}({\mathsf{P}}_{\mu,i}\boldsymbol{\rho})=\sum_{\nu=1}^{M}\sum_{j=1}^{N_{\nu}}\mathrm{tr}({\mathsf{P}}_{\nu,j}\boldsymbol{\rho})\,\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{G}}_{\nu,j})\,.$ (89)
+It needs to show that the right-hand side of this formula indeed reduces to its left-hand side. Doing some calculations finally results in the formulas
+$\sum_{\nu\neq\mu}\sum_{j=1}^{N_{\nu}}\mathrm{tr}({\mathsf{P}}_{\nu,j}\boldsymbol{\rho})\,\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{G}}_{\nu,j})$ $=\frac{(M-1)a_{\mu}}{Md}\ ,$ (90)
+$\sum_{j=1}^{N_{\mu}}\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})\,\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{G}}_{\mu,j})$ $=\mathrm{tr}({\mathsf{P}}_{\mu,i}\boldsymbol{\rho})-\frac{(M-1)a_{\mu}}{Md}\ .$ (91)
+Combining (90) with (91) proves (89).
+Due to definition of $\boldsymbol{\varpi}$, we see from (87) that
+$\mathrm{tr}(\boldsymbol{\rho}^{2})\geq\sum_{\mu=1}^{M}\sum_{\nu=1}^{M}\sum_{i=1}^{N_{\mu}}\sum_{j=1}^{N_{\nu}}\mathrm{tr}({\mathsf{P}}_{\mu,i}\boldsymbol{\rho})\,\mathrm{tr}({\mathsf{P}}_{\nu,j}\boldsymbol{\rho})\,\mathrm{tr}\bigl{(}{\mathsf{G}}_{\mu,i}{\mathsf{G}}_{\nu,j}\bigr{)}\,.$ (92)
+It also follows from (88) that
+$\mathrm{tr}\bigl{(}{\mathsf{G}}_{\mu,i}{\mathsf{G}}_{\nu,j}\bigr{)}=\frac{\delta_{\mu\nu}\delta_{ij}a_{\mu}^{2}(b_{\mu}-c_{\mu})+\delta_{\mu\nu}a_{\mu}^{2}(c_{\mu}-f)}{a_{\mu}^{2}(b_{\mu}-c_{\mu})a_{\nu}^{2}(b_{\nu}-c_{\nu})}+\frac{f}{M^{2}\gamma_{\mu}\gamma_{\nu}}\ .$ (93)
+Substituting (93) into the right-hand side of (92) leads to the different sums. By calculations, the sum with the last term of (93) reads as
+$\frac{f}{M^{2}}\sum_{\mu=1}^{M}\,\frac{1}{\gamma_{\mu}}\sum_{i=1}^{N_{\mu}}\mathrm{tr}({\mathsf{P}}_{\mu,i}\boldsymbol{\rho})\sum_{\nu=1}^{M}\,\frac{1}{\gamma_{\nu}}\sum_{j=1}^{N_{\nu}}\mathrm{tr}({\mathsf{P}}_{\nu,j}\boldsymbol{\rho})=\frac{f}{M^{2}}\sum_{\mu=1}^{M}\sum_{\nu=1}^{M}1=f\,.$ (94)
+Substituting the first fraction in the right-hand side of (93) into (92) allows one to rewrite this inequality as
+$\mathrm{tr}(\boldsymbol{\rho}^{2})-f\geq\sum_{\mu=1}^{M}\sum_{j=1}^{N_{\mu}}\frac{\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})^{2}}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}+\sum_{\mu=1}^{M}\frac{\gamma_{\mu}^{2}(c_{\mu}-f)}{a_{\mu}^{2}(b_{\mu}-c_{\mu})^{2}}=\sum_{\mu=1}^{M}\sum_{j=1}^{N_{\mu}}\frac{\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})^{2}}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}-\frac{1}{d}\,\sum_{\mu=1}^{M}\frac{\gamma_{\mu}}{a_{\mu}(b_{\mu}-c_{\mu})}\ .$ (95)
+Here, we omitted some details based on the use of (8). It is clear that (95) is equivalent to the aim (86). $\blacksquare$
+The inequality (85) allows us to estimate the index of coincidence averaged with the weights (34).
+In special cases, it reduces to the results obtained previously, including a set of MUBs [70], MUMs [51] and $(N,M)$-POVMs [23].
+In all these cases, the weights (34) turned out to be equal.
+We refrain from presenting the details here.